# Learning Penalty for Optimal Partitioning via Automatic Feature Extraction

### v1

@@ 3 Novelty
In a Recurrent Network with input size $k\geq 1$ and hidden size $m\geq 1$, processing a sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]\in\mathbb{R}^{N}$, the hidden state at time $t$, called $h_{t}$, is updated as:
$h_{t}=f(d_{[t:t+k-1]};h_{t-1})\in\mathbb{R}^{m}$
where $d_{[t:t+k-1]}=[d_{t},d_{t+1},...,d_{t+k-1}]\in\mathbb{R}^{k}$. Time stamp $t$ ranges from $k$ to $N$. The initial hidden state $h_{k-1}$ can be assigned manually. The final hidden state $h_{N}\in\mathbb{R}^{m}$ captures information from the entire sequence. Thus, we use $h_{N}$ as the $m$ extracted features from the raw sequence.
Predicting $\lambda$ from extracted features.
The prediction is formed by the linear combination of $m$ extracted features:
$\log\lambda=h_{N}\cdot\beta+\beta_{0}$
@@
Figure 5: Example: visualization of features vs. targets from the dataset detailed. The upper plots show the relationship between length or variance and the target, as these features are considered key based on $\lambda$ prediction unsupervised methods BIC and AIC. The two lower plots illustrate the visualization of automatically extracted features vs. targets, using a GRU model with one hidden layer and a hidden size of 2, which corresponds to the number of extracted features. Since we consider that the predicted $\lambda$ has a linear relationship with the features, a good feature should exhibit an approximately linear relationship in the plot. In other words, a good feature should allow us to intuitively draw a straight line that above the lower targets (blue dots) and below the upper targets (red dots). From the figure, it is evident that Extracted Feature 1 is a better feature (compared to length or variance), as we can visually draw a clear line that separates the lower and upper targets.
Automatic extracted features vs. Manual extracted features.
The novel method generates useful features without the need for a large number, unlike previous models. Figure 5 provides an example by comparing the ability to predict $\lambda$ using length (or variance) - as these manual extracted features are considered key for $\lambda$ prediction in unsupervised methods - versus automatic extracted features from recurrent networks. It demonstrates that the automatic extracted features probably provide a more accurate prediction of $\lambda$ than either length or variance.
@@
Recurrent networks have the ability to approximate statistical sequence features. For example, if the core network is defined as $h_{t}=h_{t-1}+1$ when $h_{0}=0$ and the input size $k=1$, then $h_{N}=N$, which represents the sequence length, a feature used in BIC. The core component of these models is a deep neural network. For instance, the Vanilla Recurrent Network (Rumelhart et al., 1986) core is a MLP, which, according to the universal approximation theorem (Hornik et al., 1989), can approximate any continuous function, regardless of its complexity, provided there are enough neurons in the hidden layer. This is why, although these models may not directly produce exact statistical sequence features, their architecture enables them to approximate these features with a high degree of accuracy.
Recurrent networks to be implemented.
This study explores three recurrent networks: Vanilla Recurrent Network (RNN), Long Short-Term Memory (LSTM) (Hochreiter, 1997), Gated Recurrent Unit (GRU) (Chung et al., 2014). RNNs struggle with long-term dependencies due to the vanishing gradient problem. LSTMs address this by introducing memory cells and gating mechanisms that help retain information over longer sequences. GRUs, a simplified version of LSTMs, combine the forget and input gates into a single update gate, offering a more efficient alternative while still capturing long-term dependencies.
4 Experiments
In this section, the implementation of models is detailed to enhance reproducibility. Labeled sequences are first processed, with features manually extracted to replicate previous work, along with the corresponding target intervals. Baseline models and the proposed models are then implemented. For each test sequence, OPART is applied with the predicted $\lambda$ to detect changepoints. Finally, accuracy rates are calculated based on the predicted changepoints and the test set labels.
@@ 4 Experiments
Except for the three datasets: systematic, detailed, and cancer, each ChIP-seq dataset is named using the following format: for example, ATAC_JV_adipose. The first part, “ATAC”, indicates the assay type. The second part, “JV”, represents the initials of the biologist who assigned the labels. The last part, “adipose”, refers to the sample set.
@@
This study uses 20 datasets: two DNA copy number profiles from neuroblastoma tumors (Rigaill et al., 2013), known for detailed and systematic dataset; one dataset cancer from (Hocking and Srivastava, 2023); and 17 ChIP-seq datasets from (Hocking et al., 2016), see Table 1 for all benchmark datasets.
@@
The evaluation metric used is the accuracy rate on the test set, using the label errors number and the total labels number.
@@ 4.1 Model Implementation
Model Regularization Citation
linear L1 Rigaill et al. (2013)
MMIT
max depth
min split sample
loss type
loss margin
Drouin et al. (2017)
AFT_XGboost
learning rate
max depth
min child weight
reg alpha
reg lambda
Barnwal et al. (2022)
constant none Baseline
@@ 4.1.1 Baseline Models
R package penaltyLearning (Hocking, 2024) was used to implement the Max Margin Interval Regression model with L1 regularization. The regularization parameter L1 starts at 0.001, increasing by a factor of 1.2 until no features remain, with cross-validation determining the optimal L1 regularization value.
@@
The MMIT model was implemented using the R package mmit (Drouin, 2017). Cross-validation is used to select optimal hyperparameters, including:
  • 1.
@@
  • 3.
    margin ($\epsilon$ in Hinge Error 1): values of 0, 1, and 2
  • 4.
    loss_type: values of hinge and square (equivalent to $p=1$ and $p=2$ in Hinge Error 1);
@@
The model is implemented using the Python package xgboost. Following the same setup described in Barnwal et al. (2022), cross-validation was performed to select hyperparameters, including:
  • 1.
@@
This model predicts a single value $\lambda^{*}$ from the train set using only target intervals. The objective is to minimize the total Hinge Error over $n$ instances with target intervals $[\lambda_{l}^{i},\lambda_{u}^{i}]_{i=1}^{n}$, defined as:
$\mathcal{L}(\lambda)=\sum_{i=1}^{n}l(\lambda,[\lambda_{l}^{i},\lambda_{u}^{i}]),\quad\lambda^{*}=\arg\min_{\lambda}\mathcal{L}(\lambda).$
@@ 4.2 Results
Figure 6 presents the accuracy rates for three datasets using the original, unpooled sequences, while Figure 7 shows the results for various ChIP-seq datasets (pooled sequences). Each dataset is evaluated using seven different methods, producing multiple accuracy values depending on the number of folds used for that dataset.
@@ 5 Discussion and Conclusion
Recurrent networks, in general, outperform previous models across datasets, see Figure 6 and Figure 7. However, for smaller datasets (with fewer than 100 instances), recurrent networks and MLPs do not consistently outperform baseline models. On the other hand, when dealing with larger datasets, recurrent networks noticeably outperform previous models and show a slight edge over MLPs.
@@
For these particular genome datasets, it is assumed that $\lambda$ exhibits an increasingly monotonic behavior with respect to both sequence length and variance, which are employed in unsupervised methods. One approach is to leverage monotonic network architectures, such as min-max networks (Sill, 1997), or more advanced architectures like Lattice Networks (You et al., 2017), with modifications to the loss function for each linear region.
We can consider CNNs (LeCun et al., 1998) because they are capable of handling sequences of varying lengths and generating a single output by utilizing convolution layers to extract features across the sequence. Global pooling layers (e.g., Global Max or Average Pooling) aggregate these features into a fixed-size representation. This vector is then passed through a MLP to produce the final output.
@@
  • 2.
    Attention mechanisms (Bahdanau et al., 2014) dynamically focus on relevant input parts, assigning importance weights to subsequences. Unlike sequential RNNs, they effectively capture long-range dependencies and handle varying input lengths, making them ideal for this task.
  • 3.
    Bidirectional Recurrent Networks (Schuster and Paliwal, 1997): The models implemented in this study are unidirectional, leading to challenges with long-term memory, since the later parts of the sequence tend to be more important than the earlier ones. Using a bidirectional recurrent network architecture can aim to address this limitation, enabling the model to better capture information from the sequence.
Declaration of generative AI technologies in the writing process
The author(s) utilized OpenAI’s ChatGPT to help rephrase sentences for improved clarity during the preparation of this work. All content was subsequently reviewed and edited by the author(s), who take full responsibility for the final published material.
@@
Appendix A Optimal Partitioning Optimization Problem
Given a sequence $\mathbf{d}\in\mathbb{R}^{N}$ and penalty parameter $\lambda\geq 0$, find the partition vector $\mathbf{m}\in\mathbb{R}^{N}$ that minimizes the following cost function:
$C_{\lambda}(\textbf{d},\textbf{m})=\sum_{i=1}^{N}l(d_{i},m_{i})+\lambda\sum_{i=1}^{N-1}I[m_{i}\neq m_{i+1}]$
Here, $l(m_{i},d_{i})$ represents the squared error between the sequence value $d_{i}$ and the mean segment $m_{i}$, i.e., $l(d_{i},m_{i})=(m_{i}-d_{i})^{2}$. The indicator function $I[m_{i}\neq m_{i+1}]$ equals 1 if a changepoint occurs ($m_{i}\neq m_{i+1}$) and 0 otherwise. Therefore, $\sum_{i=1}^{N-1}I[m_{i}\neq m_{i+1}]$ denotes the total number of changepoints.
Appendix B Statistics on the accuracy of each dataset with respect to each method
@@
Dataset constant linear mmit AFT_XGboost mlp rnn lstm gru

### v2

@@ 3 Novelty
In a Recurrent Network with input size $k\geq 1$ and hidden size $m\geq 1$, processing a sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]\in\mathbb{R}^{N}$, the hidden state at time $t$, called $h_{t}$, using a function $f$, is updated as:
$h_{t}=f(d_{[t-k+1:t]};h_{t-1})\in\mathbb{R}^{m}$ (4)
where $d_{[t-k+1:t]}=[d_{t-k+1},d_{t-k+2},...,d_{t}]\in\mathbb{R}^{k}$. Time stamp $t$ ranges from $k$ to $N$. The initial hidden state $h_{k-1}$ can be assigned manually. The final hidden state $h_{N}\in\mathbb{R}^{m}$ captures information from the entire sequence. Thus, we use $h_{N}$ as the $m$ automatically extracted features from the raw sequence.
$\lambda$ prediction.
The prediction is the linear combination of $m$ extracted features:
$\log\hat{\lambda}=h_{N}\cdot\beta+\beta_{0}\text{\quad or \quad}\hat{\lambda}=e^{h_{N}\cdot\beta+\beta_{0}}$
@@
Recurrent networks have the ability to approximate sequence features. The core component of these models is a neural network. For instance, the Vanilla Recurrent Network [33] core is a fully connected neural network, which, according to the universal approximation theorem [34], can approximate any continuous function, regardless of its complexity, provided there are enough neurons in the hidden layer. This is why, although these models may not directly produce exact statistical sequence features, their architecture enables them to approximate these features with a high degree of accuracy. Refer to Table 1 to see how recurrent networks are capable of approximating sequence features.
Figure 4: Example: features vs. targets from the dataset detailed. The upper plots: the length or variance vs. targets, as these features are used on $\lambda$ prediction unsupervised methods. The lower plots: automatically extracted features vs targets, using a GRU model with two extracted features. Since we consider the predicted $\lambda$ has a linear relationship with the features, a good feature should exhibit a fairly linear pattern in the plot.
Automatic vs. Manual extracted features.
Figure 4 provides an example by comparing the ability to predict $\lambda$ using length (or variance) - as these features are considered key for $\lambda$ prediction in unsupervised methods - versus automatic extracted features from recurrent networks. Since we assume that the predicted value of $\lambda$ has a linear relationship with the features, a good feature should display a linear pattern in the plot. It can be seen that the automatically extracted Feature 1 exhibits an excellent linear pattern with the all the target intervals since it’s easy to visualize a line intersecting most of these intervals.
Recurrent network architectures.
This study explores 3 recurrent networks: Vanilla Recurrent Network (RNN), Long Short-Term Memory (LSTM) [35], Gated Recurrent Unit (GRU) [36]. RNNs struggle with long-term dependencies due to the vanishing gradient problem. LSTMs address this by using memory cells and gating mechanisms that help retain information over longer sequences. GRUs, a simplified version of LSTMs, combine the forget and input gates into a single update gate, offering a more efficient alternative while still capturing long-term dependencies.
@@
4 Experiments
In this section, the implementation of models is detailed. Labeled sequences are first processed, with features manually extracted to replicate previous work, along with the corresponding target intervals. Baseline models and the proposed models are then implemented. For each test sequence, OPART is applied with the predicted $\lambda$ to detect changepoints. Finally, accuracy rates are calculated based on the predicted changepoints and the test set labels.
@@ 4 Experiments
This study uses 20 datasets: two DNA copy number profiles from neuroblastoma tumors [19], known for detailed and systematic dataset; one dataset cancer from [15]; and 17 ChIP-seq datasets from [37], see Table 2 for all benchmark datasets, each sequence has a different length (number of datapoints), so the range of sequence lengths within each dataset is reported in the Table 2. For each sequence in each dataset, biologist experts use their visual interpretation of important signal and noise patterns to label regions as having changes in mean value or not.
@@
The metric used is the accuracy rate on the test set:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
where
  • TP = True Positives (The number of detected changepoints matches to the positive labels)
  • TN = True Negatives (No detected changepoints in the negative labels)
  • FP = False Positives (The number of detected changepoints exceeds that of the labels.)
  • FN = False Negatives (The number of detected changepoints is less than that of the labels.)
@@ 4.1 Model Implementation
Model Regularization Language Package Paper
linear L1 R [38] [19]
MMIT tree architecture Python/Cpp binding [39] [20]
AFT_XGboost tree architecture Python [40] [21]
constant none Python/Cpp binding [39] Baseline
@@ 4.1.1 Baseline Models
This model uses L1 regularization. The regularization parameter L1 starts at 0.001, geometrically increasing by a factor of 1.2 until no features remain, with cross-validation determining the optimal L1 regularization value.
@@
The optimal hyperparameters are chosen through cross-validation on the train training set, including:
@@
  • margin ($\epsilon$ in Hinge Error 3): values of 0, 1, and 2
  • loss_type: values of hinge and square (equivalent to $p=1$ and $p=2$ in Hinge Error 3);
@@
Following the same setup described in Barnwal et al. [21], the optimal hyperparameters are chosen through cross-validation on the train training set, including:
@@
This model predicts a single value $\hat{\lambda}=\exp{\log\hat{\lambda}}$ from the train set using only target intervals. The objective is to minimize the total Hinge Error (3) over $S$ instances with target intervals $\log\lambda_{l}^{i},\log\lambda_{u}^{i}]_{i=1}^{S}$:
$\log\hat{\lambda}=\operatorname*{argmin}_{c}\sum_{i=1}^{S}l(c,[\log\lambda_{l}^{i},\log\lambda_{u}^{i}]).$
@@ 4.2 Results
The comprehensive accuracy results of all 8 methods for each of 20 datasets are provided in the Appendix, as the table is too large to be included in the main paper. Figure 5 presents the accuracy rates for three datasets using the original, unpooled sequences, while Figure 6 shows the results for various ChIP-seq datasets (pooled sequences). Each dataset is evaluated using seven different methods, producing multiple accuracy values depending on the number of folds used for that dataset.
@@ 5 Discussion and Conclusion
Recurrent networks, in general, outperform previous models across datasets, see Figure 5 and Figure 6. For small datasets (with fewer than 100 instances), the accuracy of the proposed models is comparable (neither better nor worse) to that of existing models. On the other hand, when dealing with larger datasets, recurrent networks noticeably outperform previous models and show a slight edge over MLPs.
@@
Practical Considerations in Real-World Applications.
There is a trade-off between training time and accuracy. The proposed models have higher accuracy in most of the 20 datasets, but they are significantly more time-consuming to train, especially for longer sequences. In contrast, manual extracted features models do not face this issue. Therefore, in real-world scenarios where datasets contain only a few sequences but each sequence is long, and training time is a priority, manual extracted features models may be the better choice for two reasons: first, their training time is significantly faster; and second, these models’ accuracy is comparable to the proposed models, which can be seen in the bottom three panels of Figure 6. However, if the user’s priority is accuracy and the hardware requirements are met, the proposed models are a better choice. Although the training process is time-consuming, it only needs to be done once; afterwards, the trained models can be applied to unseen sequences, as their feed-forward process is sufficiently fast.
@@
For these particular datasets, it is assumed that $\lambda$ exhibits an increasingly monotonic behavior with respect to both sequence length and variance, which are employed in unsupervised methods. One approach is to leverage monotonic network architectures, such as min-max networks [41], or more advanced architectures like Lattice Networks [42], with modifications to the loss function for each linear region.
We can consider CNNs [43] because they are capable of handling sequences of varying lengths and generating a single output by utilizing convolution layers to extract features across the sequence. Global pooling layers (e.g., Global Max or Average Pooling) aggregate these features into a fixed-size representation. This vector is then passed through a MLP to produce the final output.
@@
  • Attention mechanisms [44] dynamically focus on relevant input parts, assigning importance weights to subsequences. Unlike sequential RNNs, they effectively capture long-range dependencies and handle varying input lengths, making them ideal for this task.
  • Bidirectional Recurrent Networks [45]: The models implemented in this study are unidirectional, leading to challenges with long-term memory, since the later parts of the sequence tend to be more important than the earlier ones. Using a bidirectional recurrent network architecture can aim to address this limitation, enabling the model to better capture information from the sequence.
@@
Appendix
@@
Dataset constant linear mmit AFT_XGB mlp rnn lstm gru